# NLP-ADBench: NLP Anomaly Detection Benchmark

### v1

@@ Appendix A Details on NLP-ADBench
A.1. Additional Details on Benchmark Datasets
A.1.1. Datasets Sources.
  1. (1)
@@ A.1. Additional Details on Benchmark Datasets
    NLPAD-AGNews is constructed from the AG News dataset (Rai, 2023), which was originally intended for news topic classification tasks. The AG News dataset contains 127,600 samples categorized into four classes: World, Sports, Business, and Sci/Tech. We selected the text from the “description” column as NLPAD-AGNews’s text data source. The “World” category was designated as the anomaly class and was downsampled accordingly.
  2. (2)
@@
    NLPAD-BBCNews is constructed from the BBC News dataset (Greene and Cunningham, 2006), which was originally used for document classification across various news topics. The BBC News dataset includes 2,225 articles divided into five categories: Business, Entertainment, Politics, Sport, and Tech. We selected the full text of the news articles as NLPAD-BBC News’s text data source. The “Entertainment” category was designated as the anomaly class and was downsampled accordingly.
  3. (3)
@@
    NLPAD-EmailSpam is constructed from the Spam Emails dataset (Metsis et al., 2006), originally used for email spam detection. The Spam Emails dataset contains 5,171 emails labeled as either spam or ham (not spam). We selected the text from the “text” column containing the email bodies as NLPAD-Emails Spam’s text data source. The “spam” category was designated as the anomaly class and was downsampled accordingly.
  4. (4)
@@
    NLPAD-Emotion: is constructed from the Emotion dataset (Saravia et al., 2018) , which was originally intended for emotion classification tasks in textual data. The Emotion dataset contains 416,809 text samples labeled with six emotions: anger, fear, joy, love, sadness, and surprise. We selected the text from the “text” column as NLPAD-Emotion’s text data source. The “fear” category was designated as the anomaly class and was downsampled accordingly.
  5. (5)
@@
    NLPAD-MovieReview: is constructed from the Movie Review dataset (Maas et al., 2011) , commonly used for sentiment analysis of film critiques. The Movie Review dataset includes 50,000 reviews labeled as positive or negative. We selected the full review texts as NLPAD-MovieReview’s text data source. The “neg” (negative reviews) category was designated as the anomaly class and was downsampled accordingly.
  6. (6)
@@
    NLPAD-N24News is constructed from the N24News dataset (Wang et al., 2022), originally used for topic classification of news articles. N24News contains 61,235 articles across various categories. We selected the full text of the news articles as NLPAD-N24News’s text data source. The “food” category was designated as the anomaly class and was downsampled accordingly.
  7. (7)
@@
    NLPAD-SMSSpam is constructed from the SMS Spam Collection dataset (Almeida et al., 2011), originally intended for classifying SMS messages as spam or ham (not spam). The SMS Spam Collection dataset comprises 5,574 messages labeled accordingly. We selected the text from the “message text” as NLPAD-SMS Spam’s text data source. The “spam” category was designated as the anomaly class and was downsampled accordingly.
  8. (8)
@@
    NLPAD-YelpReview is constructed from the Yelp Review Polarity dataset  (Putra, 2023), originally intended for sentiment classification tasks. The Yelp Review Polarity dataset is created by considering 1-star and 2-star ratings as negative, and 3-star and 4-star ratings as positive. For each polarity, 280,000 training samples and 19,000 testing samples were randomly selected, resulting in a total of 560,000 training samples and 38,000 testing samples. Negative polarity is labeled as class 1, and positive polarity as class 2. We selected the text from the text column as NLPAD-YelpReview’s text data source. The label 1 (negative reviews) was designated as the anomaly class and was downsampled accordingly.
A.1.2. NLPAD dataset’s text pre-processing
@@
On all 8 datasets, we preprocessed the raw text data to ensure consistency and usability by removing URLs and HTML tags, eliminating unnecessary special characters while retaining essential punctuation, converting line breaks and consecutive spaces into single spaces, and preserving case sensitivity and stop words to maintain linguistic integrity. After processing the text, we found that some texts became duplicates due to the removal of certain symbols. Consequently, we removed all duplicate data to ensure the uniqueness of each text sample. These preprocessing steps follow established practices to effectively clean text data while retaining its syntactic and semantic features, providing a reliable foundation for natural language processing tasks (Chai, 2022).
A.2. Additional Details on Algorithms
A.2.1. End-to-End Algorithms
  1. (1)
@@ A.2.1. End-to-End Algorithms
    Context Vector Data Description (Ruff et al., 2019)(CVDD) is an unsupervised anomaly detection method for textual data. It utilizes pre-trained word embeddings and a multi-head self-attention mechanism to learn ”context vectors” that represent normal patterns in the data. Anomalies are detected by measuring the cosine distance between sequence projections and context vectors, where larger distances indicate higher anomaly likelihoods. CVDD also penalizes overlapping contexts to enhance interpretability.
  2. (2)
@@
    Detecting Anomalies in Text via Self-Supervision of Transformers (DATE) (Manolache et al., 2021) detects anomalies in text by training self-supervised transformers on tasks like replaced mask detection, enabling the model to learn normal language patterns and identify deviations.
  3. (3)
@@
    Few-shot Anomaly Detection in Text with Deviation Learning (FATE) (Das et al., 2023) is a deep learning framework that uses a small number of labeled anomalies to learn anomaly scores end-to-end. By employing deviation learning, it ensures normal examples align with reference scores while anomalies deviate significantly. Utilizing multi-head self-attention and multiple instance learning, FATE achieves state-of-the-art performance on benchmark datasets.
A.2.2. Traditional Algorithms
  1. (1)
@@ A.2.2. Traditional Algorithms
    Local Outlier Factor (LOF) (Breunig et al., 2000) calculates the local density deviation of a data point relative to its neighbors. This metric identifies points that have substantially lower density than their neighbors, marking them as outliers.
  2. (2)
@@
    Deep Support Vector Data Description (DeepSVDD) (Ruff et al., 2018) minimizes the volume of a hypersphere enclosing the data representations learned by a neural network, capturing common patterns while identifying anomalies as points outside the hypersphere.
  3. (3)
@@
    Empirical-Cumulative-distribution-based Outlier Detection (ECOD)  (Li et al., 2022) estimates the empirical cumulative distribution function (ECDF) for each feature independently. It identifies outliers as data points that reside in the tails of these distributions. This approach is hyperparameter-free and offers straightforward interpretability.
  4. (4)
@@
    Isolation Forest (IForest) (Liu et al., 2008) detects anomalies by isolating observations through random feature selection and splitting, with anomalies requiring fewer splits
  5. (5)
@@
    Single-Objective Generative Adversarial Active Learning (SO_GAAL) (Liu et al., 2019) optimizes a single objective function to generate adversarial samples and effectively identify anomalies in unsupervised settings.
  6. (6)
@@
    AutoEncoder (AE) (Aggarwal, 2017) detects anomalies by reconstructing input data, where higher reconstruction errors signify potential anomalies.
  7. (7)
@@
    Unifying Local Outlier Detection Methods via Graph Neural Networks(LUNAR) (Goodge et al., 2022) uses graph neural networks to integrate and enhance traditional local outlier detection methods, unifying them for better anomaly detection.
  8. (8)
@@
    Variational AutoEncoder (VAE) (Kingma and Welling, 2013; Burgess et al., 2018) uses probabilistic latent variables to model data distributions, identifying anomalies based on reconstruction probabilities or latent space deviations.

### v2

@@ 4 Conclusion
Ethics Statement
This work adheres to ethical standards emphasizing transparency, fairness, and privacy in NLP anomaly detection research. By openly sharing datasets, algorithms, and experimental results, NLP-ADBench provides a standardized foundation for advancing safer and more reliable web-based systems. All datasets are publicly available and contain no personally identifiable information, ensuring privacy compliance. Pre-trained embeddings (such as OpenAI’s text-embedding-3-large) are used in accordance with their terms of service. Additionally, we used ChatGPT exclusively to improve minor grammar in the final manuscript text.
Broader Impacts
The NLP-ADBench proposed in this paper provides a comprehensive benchmark framework for anomaly detection in NLP. By standardizing datasets and algorithms, this work supports advancements in critical web-based applications, including fraud detection, spam filtering, and content moderation. The benchmark promotes transparency, reproducibility, and facilitates further innovations, ultimately contributing to safer, more reliable online environments.
Acknowledgments
This work was partially supported by the National Science Foundation under Award Nos. 2428039, 2346158, and 2449280. We also acknowledge the use of computational resources provided by the Advanced Cyberinfrastructure Coordination Ecosystem Boerner et al. (2023): Services & Support (ACCESS) program, supported by NSF grants #2138259, #2138286, #2138307, #2137603, and #2138296. Specifically, this work used NCSA Delta GPU at the National Center for Supercomputing Applications (NCSA) through allocation CIS250073. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. The authors also gratefully acknowledge support from the Amazon Research Awards and Capital One Research Awards.
@@ Appendix A Details on NLP-ADBench
A.1 Additional Details on Benchmark Datasets
A.1.1 Datasets Sources.
  1. 1.
@@ A.1 Additional Details on Benchmark Datasets
    NLPAD-AGNews is constructed from the AG News dataset Rai (2023), which was originally intended for news topic classification tasks. The AG News dataset contains 127,600 samples categorized into four classes: World, Sports, Business, and Sci/Tech. We selected the text from the “description” column as NLPAD-AGNews’s text data source. The “World” category was designated as the anomaly class and was downsampled accordingly.
  2. 2.
@@
    NLPAD-BBCNews is constructed from the BBC News dataset Greene and Cunningham (2006), which was originally used for document classification across various news topics. The BBC News dataset includes 2,225 articles divided into five categories: Business, Entertainment, Politics, Sport, and Tech. We selected the full text of the news articles as NLPAD-BBC News’s text data source. The “Entertainment” category was designated as the anomaly class and was downsampled accordingly.
  3. 3.
@@
    NLPAD-EmailSpam is constructed from the Spam Emails dataset Metsis et al. (2006), originally used for email spam detection. The Spam Emails dataset contains 5,171 emails labeled as either spam or ham (not spam). We selected the text from the “text” column containing the email bodies as NLPAD-Emails Spam’s text data source. The “spam” category was designated as the anomaly class and was downsampled accordingly.
  4. 4.
@@
    NLPAD-Emotion: is constructed from the Emotion dataset Saravia et al. (2018) , which was originally intended for emotion classification tasks in textual data. The Emotion dataset contains 416,809 text samples labeled with six emotions: anger, fear, joy, love, sadness, and surprise. We selected the text from the “text” column as NLPAD-Emotion’s text data source. The “fear” category was designated as the anomaly class and was downsampled accordingly.
  5. 5.
@@
    NLPAD-MovieReview: is constructed from the Movie Review dataset Maas et al. (2011) , commonly used for sentiment analysis of film critiques. The Movie Review dataset includes 50,000 reviews labeled as positive or negative. We selected the full review texts as NLPAD-MovieReview’s text data source. The “neg” (negative reviews) category was designated as the anomaly class and was downsampled accordingly.
  6. 6.
@@
    NLPAD-N24News is constructed from the N24News dataset Wang et al. (2022), originally used for topic classification of news articles. N24News contains 61,235 articles across various categories. We selected the full text of the news articles as NLPAD-N24News’s text data source. The “food” category was designated as the anomaly class and was downsampled accordingly.
  7. 7.
@@
    NLPAD-SMSSpam is constructed from the SMS Spam Collection dataset Almeida et al. (2011), originally intended for classifying SMS messages as spam or ham (not spam). The SMS Spam Collection dataset comprises 5,574 messages labeled accordingly. We selected the text from the “message text” as NLPAD-SMS Spam’s text data source. The “spam” category was designated as the anomaly class and was downsampled accordingly.
  8. 8.
@@
    NLPAD-YelpReview is constructed from the Yelp Review Polarity dataset  Putra (2023), originally intended for sentiment classification tasks. The Yelp Review Polarity dataset is created by considering 1-star and 2-star ratings as negative, and 3-star and 4-star ratings as positive. For each polarity, 280,000 training samples and 19,000 testing samples were randomly selected, resulting in a total of 560,000 training samples and 38,000 testing samples. Negative polarity is labeled as class 1, and positive polarity as class 2. We selected the text from the text column as NLPAD-YelpReview’s text data source. The label 1 (negative reviews) was designated as the anomaly class and was downsampled accordingly.
A.1.2 NLPAD dataset’s text pre-processing
@@
On all 8 datasets, we preprocessed the raw text data to ensure consistency and usability by removing URLs and HTML tags, eliminating unnecessary special characters while retaining essential punctuation, converting line breaks and consecutive spaces into single spaces, and preserving case sensitivity and stop words to maintain linguistic integrity. After processing the text, we found that some texts became duplicates due to the removal of certain symbols. Consequently, we removed all duplicate data to ensure the uniqueness of each text sample. These preprocessing steps follow established practices to effectively clean text data while retaining its syntactic and semantic features, providing a reliable foundation for natural language processing tasks Chai (2022).
A.2 Additional Details on Algorithms
A.2.1 End-to-End Algorithms
  1. 1.
@@ A.2.1 End-to-End Algorithms
    Context Vector Data Description Ruff et al. (2019)(CVDD) is an unsupervised anomaly detection method for textual data. It utilizes pre-trained word embeddings and a multi-head self-attention mechanism to learn "context vectors" that represent normal patterns in the data. Anomalies are detected by measuring the cosine distance between sequence projections and context vectors, where larger distances indicate higher anomaly likelihoods. CVDD also penalizes overlapping contexts to enhance interpretability.
  2. 2.
@@
    Detecting Anomalies in Text via Self-Supervision of Transformers (DATE) Manolache et al. (2021) detects anomalies in text by training self-supervised transformers on tasks like replaced mask detection, enabling the model to learn normal language patterns and identify deviations.
  3. 3.
@@
    Few-shot Anomaly Detection in Text with Deviation Learning (FATE) Das et al. (2023) is a deep learning framework that uses a small number of labeled anomalies to learn anomaly scores end-to-end. By employing deviation learning, it ensures normal examples align with reference scores while anomalies deviate significantly. Utilizing multi-head self-attention and multiple instance learning, FATE achieves state-of-the-art performance on benchmark datasets. However, as our approach focuses on unsupervised anomaly detection, we adapt FATE into FATE* by training exclusively on normal data. This adaptation involves modifying the framework to learn reference scores and deviations without access to labeled anomalies, enabling effective detection of anomalous examples in an entirely unsupervised setting.
A.2.2 Traditional Algorithms
  1. 1.
@@ A.2.2 Traditional Algorithms
    Local Outlier Factor (LOF) Breunig et al. (2000) calculates the local density deviation of a data point relative to its neighbors. This metric identifies points that have substantially lower density than their neighbors, marking them as outliers.
  2. 2.
@@
    Deep Support Vector Data Description (DeepSVDD) Ruff et al. (2018) minimizes the volume of a hypersphere enclosing the data representations learned by a neural network, capturing common patterns while identifying anomalies as points outside the hypersphere.
  3. 3.
@@
    Empirical-Cumulative-distribution-based Outlier Detection (ECOD)  Li et al. (2022) estimates the empirical cumulative distribution function (ECDF) for each feature independently. It identifies outliers as data points that reside in the tails of these distributions. This approach is hyperparameter-free and offers straightforward interpretability.
  4. 4.
@@
    Isolation Forest (IForest) Liu et al. (2008) detects anomalies by isolating observations through random feature selection and splitting, with anomalies requiring fewer splits
  5. 5.
@@
    Single-Objective Generative Adversarial Active Learning (SO_GAAL) Liu et al. (2019) optimizes a single objective function to generate adversarial samples and effectively identify anomalies in unsupervised settings.
  6. 6.
@@
    AutoEncoder (AE) Aggarwal (2017) detects anomalies by reconstructing input data, where higher reconstruction errors signify potential anomalies.
  7. 7.
@@
    Unifying Local Outlier Detection Methods via Graph Neural Networks(LUNAR) Goodge et al. (2022) uses graph neural networks to integrate and enhance traditional local outlier detection methods, unifying them for better anomaly detection.
  8. 8.
@@
    Variational AutoEncoder (VAE) Kingma and Welling (2013); Burgess et al. (2018) uses probabilistic latent variables to model data distributions, identifying anomalies based on reconstruction probabilities or latent space deviations.
A.3 More Experiment Results
We also report AUPRC scores (Table. A1) for all 19 algorithms across the 8 NLPAD datasets, along with their average AUPRC ranks (Fig. A1), to provide a complementary evaluation perspective beyond AUROC.
Table A1: Performance comparison of 19 Algorithms on 8 NLPAD datasets using AUPRC, with best results highlighted in bold and shaded.
CVDD 0.1296 0.2976 0.5353 0.0955 0.1576 0.2886 0.0712 0.1711
DATE 0.3996 0.5764 0.8885 0.1619 0.1682 0.2794 0.6112 0.2149
FATE* 0.2787 0.5805 0.5529 0.1026 0.1752 0.2777 0.1257 0.2112
BERT + LOF 0.2549 0.6029 0.2370 0.1170 0.1621 0.1678 0.1837 0.2629
BERT + DeepSVDD 0.2160 0.1328 0.2117 0.0986 0.1387 0.0798 0.1178 0.2174
BERT + ECOD 0.1616 0.2037 0.2077 0.1024 0.1374 0.0928 0.1156 0.2197
BERT + iForest 0.1559 0.2131 0.1894 0.1007 0.1412 0.0872 0.0994 0.2203
BERT + SO-GAAL 0.1033 0.0849 0.1130 0.1036 0.1486 0.0837 0.0714 0.2440
BERT + AE 0.2232 0.4274 0.2937 0.1037 0.1479 0.1255 0.1914 0.2525
BERT + VAE 0.1878 0.2559 0.2247 0.1019 0.1405 0.0957 0.1360 0.2331
BERT + LUNAR 0.2717 0.5943 0.3571 0.1053 0.1497 0.1436 0.1817 0.2609
OpenAI + LOF 0.5443 0.7714 0.5967 0.2290 0.2133 0.2248 0.2450 0.5710
OpenAI + DeepSVDD 0.1062 0.1288 0.1195 0.1040 0.3278 0.1297 0.0721 0.1893
OpenAI + ECOD 0.3294 0.2424 0.5597 0.7443 0.5165 0.2238 0.0821 0.8639
OpenAI + iForest 0.1278 0.1376 0.3283 0.1311 0.1724 0.0913 0.0772 0.2527
OpenAI + SO-GAAL 0.1538 0.0665 0.1096 0.1291 0.3005 0.0963 0.1213 0.2735
OpenAI + AE 0.4022 0.7485 0.5580 0.8355 0.1969 0.1984 0.1030 0.7063
OpenAI + VAE 0.3659 0.2424 0.5604 0.7744 0.1486 0.2537 0.0812 0.8467
OpenAI + LUNAR 0.6918 0.8653 0.5810 0.3112 0.2193 0.4425 0.1640 0.4524
Figure A1: Average rank on AUPRC of 19 NLPAD methods across 8 datasets (the lower the better).